# A Simple First-Order Algorithm for Full-Rank Equality Constrained Optimization

### v1

@@ 3.7 Combined complexity
Suppose that AS.0-AS.7 hold. Then $\frac{1}{k+1}\sum_{j=0}^{k}\Big(\|g_{T,j}\|+\|c_{j}\|\Big)\leq\frac{\kappa_{\mbox{\tiny ADSW,1}}}{\sqrt{k+1}}+\frac{\kappa_{\mbox{\tiny ADSW,2}}}{k+1}={\pazocal{O}}\left(\frac{1}{\sqrt{k+1}}\right),$ (3.26) where $\kappa_{\mbox{\tiny ADSW,1}}=\kappa_{T}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\;\;\mbox{ and }\;\;\kappa_{\mbox{\tiny ADSW,2}}=\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right).$
  • Proof.    Now consider iterations from 0 to $k$ of both types (tangential and normal) by setting $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$ (as in Lemma 3.5). We then obtain, by combining (3.18) and (3.20), that
    $\sum_{j=0}^{k}$ $\big(\|g_{T,j}\|+\|c_{j}\|\Big)=\sum_{\tau=\tau_{0}}^{\tau_{1}}\Big(\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\Big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\Big(\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\Big)$
    $\leq\kappa_{T}\sqrt{k+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$
    where we used the inequalities $\tau_{1}\leq k_{\tau_{1}}\leq k$ and $k_{\nu_{1}}\leq k$. The bound (3.26) is finally obtained by dividing both sides by $k+1$. $\Box$
@@ 4 Numerical illustration
    The algorithm appears to converge as predicted. Despite the fact that our current analysis does not cover the case where the Jacobian may lose rank, the algorithm did solve a number of instances where this occurs, sometimes finding an infeasible critical point of the constraints’ violation [^5] This occurs, for instance, at the starting point of HS61..
@@
    Its performance and reliability is (as could be anticipated) dominated by that of AdaGrad for the tangential step. Since this method is purely first-order, it often performs well, but may fail to solve ill-conditioned problems in a reasonable number of iterations. This is also the case for ADSWITCH. Over the 71 problems in our test set, it solves 44 (62%) of them within 750 iterations, 58 (81%) of them within 100000 iterations and fails on 13. While reliability is not a strong point of AdaGrad and ADSWITCH in the deterministic case, the situation is very different when noise is present, as we show below.
We illustrate a few case of satisfactory convergence in Figures 1 and 2. In these figures, one can clearly see the difference in speed of convergence between tangential steps (AdaGrad-like) and normal steps (Newton), and thus that the overall performance is dominated by that of the first-order method defining the tangential step. This is even clearer when convergence is too slow, as shown in Figure 4 where constraint violation remains very small [^6] HS50 has linear equality constraints only., while very slow convergence of the projected gradient results in a large number of iterations. The left panel of Figure 1 (ORTHREGA with $n=133$ and $m=64$) is also interesting because it shows the switching condition (2.12) in action: the algorithm first approaches an infeasible critical point of the constraints’ violation until the switching rule allows the tangential step to take over, causing the iterates to escape.
@@
Figure 5: Evolution of the two components of the optimality measure for a run of the noisy BT1 problem, with 5% (top left), 15% (top right), 25% (bottom left) and 50% (bottom right) relative Gaussian noise on the gradient
@@ 5 Conclusions and perspectives
false We have proposed a very simple first-order algorithm for solving nonlinear optimization problems with nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility. It does so without using a merit function or a filter to enforce convergence, but instead relies on the simple switching condition (2.12). The tangent steps are based on the AdaGrad method for unconstrained minimization. As is the case for AdaGrad, the objective function is never evaluated. We have analyzed its worst-case evaluation complexity and obtain global convergence rates which match the best known rates for unconstrained problems. Numerical experiments have been presented indicating that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems.
At this stage, several theoretical questions remain open for further research. These include a stochastic convergence analysis, covering the cases of rank-deficient Jacobians and/or unbounded gradients, the use of alternative minimization methods in the tangential step (such as ADAM [23], ASTR1 [19], OFFAR [20] or more standard unconstrained minimizing techniques using objective function’s values). On the practical side, further numerical experiments are clearly necessary for assessing the true potential of the new method. The handling of inequality constraints is also of obvious interest.
@@ Acknowledgement
Philippe Toint is grateful for the continued and friendly support of the Algo team at Toulouse IRIT. The authors also thank Benedetta Morini and Stefania Bellavia for pointing out a number of typos.
@@
  • Table A.1 reports the results of the noiseless runs. Columns 2 and 3 indicate the number of variables and the number of constraints. The column “#its” gives the number of iterations and the column “exitc” the termination condition (“convg” = convergence, “infeas” = infeasible critical point, “maxit” = maximum number of iterations reached).
@@
Table A.1: Results of running ADSWITCH in the abscence of noise
@@
Table A.2: Results of running ADSWITCH with 5% relative Gaussian noise
@@
Table A.3: Results of running ADSWITCH with 15% relative Gaussian noise
@@
Table A.4: Results of running ADSWITCH with 25% relative Gaussian noise
@@
Table A.5: Results of running ADSWITCH with 50% relative Gaussian noise

### v2

@@ 3.7 Combined complexity
Suppose that AS0-AS7 hold and that either the gradient is exact (i.e. $g_{T,k}=G_{T,k}$ for all $k\geq 0$) or AS.8–AS.10 hold. Then $\frac{1}{k+1}\sum_{j=0}^{k}\mathbb{E}\!\left[\|G_{T,j}\|+\|c_{j}\|\right]\leq\frac{\kappa_{\mbox{\tiny ADSW}}}{\sqrt{k+1}}={\pazocal{O}}\left(\frac{1}{\sqrt{k+1}}\right),$ (3.27) where $\kappa_{\mbox{\tiny ADSW}}=\kappa_{T}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}.$
  • Proof.    Observe first that AS.8, AS.9 and AS.10 automatically hold in the gradient is exact. Now consider iterations from 0 to $k$ of both types (tangential and normal) by setting $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$ (as in Lemma 3.5). We then obtain, by combining (3.20) and (3.22), that
    $\sum_{j=0}^{k}$ $\mathbb{E}\!\left[\|G_{T,j}\|+\|c_{j}\|\right]=\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\right]+\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\right]$
    $\leq\kappa_{T}\sqrt{k+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\sqrt{k+1},$
    where we used the inequalties $\tau_{1}\leq k_{\tau_{1}}\leq k$ and $k_{\nu_{1}}\leq k$. The bound (3.27) is finally obtained by dividing both sides by $k+1$. $\Box$
We may now use a standard probabilistic argument and show that the probability of a small optimality measure $\|G_{T,k}\|+\|c_{k}\|$ increases to one when $k$ grows.
Corollary 1
Suppose that AS.0–AS.10 hold. Then $\mathbb{P}\!\left[\min_{j\in\{0,\ldots,k\}}\left[\|G_{T,j}\|+\|c_{j}\|\right]\leq\epsilon\right]\geq 1-\delta\;\;\mbox{for}\;\;k>\frac{\kappa_{\mbox{\tiny ADSW}}^{2}}{\epsilon^{2}\delta^{2}}.$
  • Proof.    Markov’s inequality and (3.27) give that
    $\mathbb{P}\!\left[\min_{j\in\{0,\ldots,k\}}\Big(\|G_{T,j}\|+\|c_{j}\|\Big)\leq\epsilon\right]$ $\geq\mathbb{P}\!\left[\frac{1}{k+1}\sum_{j=0}^{k}\Big(\|G_{T,j}\|+\|c_{j}\|\Big)\leq\epsilon\right]$
    $\geq 1-\frac{1}{\epsilon}\mathbb{E}\!\left[\sum_{j=0}^{k}\Big(\|g_{T,j}\|+\|c_{j}\|\Big)\right]$
    $\geq 1-\frac{\kappa_{\mbox{\tiny ADSW}}}{\epsilon\sqrt{k+1}}.$
    The desired conclusion follows. $\Box$
@@ 4 Numerical illustration
    The algorithm appears to converge as predicted. Despite the fact that our current analysis does not cover the case where the Jacobian may loose rank, the algorithm did solve a number of instances where this occurs, sometimes finding an infeasible critical point of the constraints’s violation [^5] This occurs, for instance, at the starting point of HS61..
@@
    Its performance and reliability is (as could be anticipated) dominated by that of AdaGrad for the tangential step. Since this method is purely first-order, it often performs well, but may fail to solve ill-conditioned problems in a reasonable number of iterations. This is also the case for ADSWITCH. Over the 71 problems in our test set, it solves 44 of them within 750 iterations, 58 of them within 100000 iterations and fails on 13. While reliability is not a strong point of AdaGrad and ADSWITCH in the deterministic case, the situation is very different when noise is present, as we show below.
We illustrate a few case of satisfactory convergence in Figures 1 and 2. In these figures, one can clearly see the difference in speed of convergence between tangential steps (AdaGrad-like) and normal steps (Newton), and thus that the overal performance is dominated by that of the first-order method defining the tangential step. This is even clearer when convergence is too slow, as shown in Figure 4 where constraint violation remains very small [^6] HS50 has linear equality constraints only., while very slow convergence of the projected gradient results in a large number of iterations. The left panel of Figure 1 (ORTHREGA with $n=133$ and $m=64$) is also interesting because it shows the switching condition (2.12) in action: the algorithm first approaches an infeasible critical point of the constraints’s violation until the switching rule allows the tangential step to take over, causing the iterates to escape.
@@
Figure 5: Evolution of the two components of the optimality measure for a run of the the noisy BT1 problem, with 5% (top left), 15% (top right), 25% (bottom left) and 50% (bottom right) relative Gaussian noise on the gradient
@@ 5 Conclusions and perspectives
We have proposed a very simple first-order algorithm for solving potentially stochastic nonlinear optimization problems with deterministic nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility. It does so without using a merit function or a filter to enforce convergence, but instead relies on the simple switching condition (2.12). The tangent steps are based on the AdaGrad method for unconstrained minimization. As is the case for AdaGrad, the objective function is never evaluated. We have analyzed its worst-case evaluation complexity, both in the deterministic and stochastic cases, and obtain global convergence rates which match the best known rates for unconstrained problems. Numerical experiments have been presented indicating that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems.
At this stage, several theoretical questions remain open for further research. These include covering the cases of rank-deficient Jacobians and/or unbounded gradients, the use of alternative minimization methods in the tangential step (such as ADAM [21], ASTR1 [17], OFFAR [18] or more standard unconstrained minimizing techniques using objective function’s values). On the practical side, further numerical experiments are clearly necessary for assessing the true potential of the new method. The handling of inequality constraints is also of obvious interest.
@@ Acknowledgement
Philippe Toint is grateful for the continued and friendly support of the Algo team at Toulouse IRIT. Both authors thank Mehi Al-Baali and the organizers of the NAOVI 2026 conference for their kind invitation to participate.
@@
  • Table A.1 reports the results of the noiseless runs. Columns 2 and 3 indicate the number of variables and the number of constraints. The colmun “#its” gives the number of iterations and the column “exitc” the termination condition (“convg” = convergence, “infeas” = infeasible critical point, “maxit” = maximum number of iterations reached).
@@
Table A.1: Results of running ADAGEC in the abscence of noise
@@
Table A.2: Results of running ADAGEC with 5% relative Gaussian noise
@@
Table A.3: Results of running ADAGEC with 15% relative Gaussian noise
@@
Table A.4: Results of running ADAGEC with 25% relative Gaussian noise
@@
Table A.5: Results of running ADAGEC with 50% relative Gaussian noise